# Follow the Energy, Find the Path: Riemannian Metrics from Energy-Based Models

### v1

@@ 1 Introduction
In this article, we answer affirmatively by proposing to derive Riemannian metrics from pretrained Energy-Based Models (EBMs) [16, 19, 20]. EBMs are a flexible class of generative models that define an energy function $E_{\theta}$, parameterized by a neural network, assigning low energy to likely data points (i.e., $p_{\theta}(\mathbf{x})\propto\exp(-E_{\theta}(\mathbf{x}))$). We show that the energy landscape of an EBM encodes a rich geometric structure and can be leveraged to derive effective Riemannian metrics. Specifically, we introduce two novel conformal Riemannian metrics—metrics that scale the identity by a positive scalar function: ${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$ proportional to the energy itself, and ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$, proportional to the inverse unnormalized density. We evaluate both against established alternatives (${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$ [13] and ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ [11]) across datasets of increasing complexity—from toy distributions with known geodesics (see § 4.2), to rotated character images where the manifold structure is partially known (see § 4.3), and finally to high-dimensional natural images where no ground truth geometry is available (see § 4.4). We show that EBM-based metrics yield geodesics that (i) remain closer to the data manifold and (ii) better reflect its intrinsic curvature (see Fig. 1 for a visualization of the geodesics).
@@ 3.1 A primer on Riemannian geometry
[…]
@@
This parameterization satisfies the boundary conditions ($\mathbf{x}_{0,\eta}{=}\mathbf{x}_{0}$, $\mathbf{x}_{1,\eta}{=}\mathbf{x}_{1}$). In Eq. 3, $\boldsymbol{\varphi}_{\eta}$ serves as a nonlinear correction to the linear path, allowing the learned path to bend toward the data manifold. We train a single interpolant network $\boldsymbol{\varphi}_{\eta}$ over batches of random endpoint pairs so it can approximate geodesics between arbitrary points (see Algo. 1). Unlike Kapusniak et al. [24], who use full autodifferentiation to compute $\dot{\mathbf{x}}_{t,\eta}$, we opt for finite difference instead. We found this approach more stable and accurate when using a fine-time discretization.
@@ 3.2 Energy-Based Models
[…]
@@
[…]
@@
EBM can be hard to train in high-dimensional pixel space, especially because of the sampling procedure [53, 54, 55]. For complex tasks, we follow standard practice and operate in the latent space of a pretrained VAE [56], where all baselines are evaluated for fairness.
@@ 3.3 EBM-derived Riemannian Metrics
Here, we describe the EBM-derived metrics ${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$, ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$. For details on the baseline Riemannian metrics ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$,${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$, see § 4.1. To ensure a fair comparison —and following standard practice in the field [42, 57, 58]— all metrics are cast using a shared parametric form:
@@
where $\mathbf{h}(\mathbf{x})$ is a metric-specific, positive-definite function (either scalar, diagonal, or matrix), and $\alpha$, $\beta$ are calibration constants. These constants are chosen so that the metric scale to $\mathbf{I}$ on the data manifold and to $10^{3}\cdot\mathbf{I}$ in low-density regions [^2] Note that this multiplicative factor amounts to a change of unit, to ensure reasonable scaling of the lengths, but the induced geodesics are only determined by the ratio $\alpha/\beta$.. Further details about the metric calibration procedure are provided in Supp. C.1. Importantly, all EBM-derived metrics are conformal, they take the form $\lambda(\mathbf{x})\mathbf{I}$, where $\lambda$ is a scalar function. In other words, they scale the identity matrix uniformly in all directions, resulting in isotropic metrics:
@@
    Intuitively, high-energy (low-density) regions receive a larger metric, penalizing movement away from the data. Note that $E_{\theta}$ is an affine rescaling of the negative log-likelihood $-\log p_{\mathcal{D}}$.
@@ 4.1 Baseline Riemannian Metrics
    ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G}_{\text{LAND}}(\mathbf{x})}=(\alpha\,\textnormal{diag}(\mathbf{h}(\mathbf{x}))+\beta\mathbf{I})^{-1}\,\,\textnormal{s.t}\,\,h^{(j)}(\mathbf{x})=\sum_{i=1}^{N}(x_{i}^{(j)}\!-\!x^{(j)})^{2}\exp{\bigg{(}\!\!\!-\frac{||\mathbf{x}\!-\!\mathbf{x_{i}}||^{2}}{2\sigma^{2}}\bigg{)}}$ (8)
@@
We evaluate ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$, ${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$, ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$, and ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ on three datasets of increasing complexity. Circular Mixture of Gaussians offers full control and ground-truth geodesics. The rotated characters dataset is higher-dimensional but still allows quantitative evaluation. Animal Faces is made of higher-dimensional images but with no ground truth. This progression tests metric performance as data complexity grows. The code to reproduce all our experiments will be released upon acceptance.
@@ 4.2 Circular Mixture of Gaussians
Metric $p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$ ()($\uparrow$)( ↑ ) RMSE ()($\downarrow$)( ↓ )
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.79$
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.77$
[…]
Metric $p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$ ()($\uparrow$)( ↑ ) RMSE ()($\downarrow$)( ↓ )
$\boldsymbol{G}_{E_{\mathcal{M}}}$ 0.67 -
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ 0.73 -
[…]
Figure 2: Geodesics on UCG and WCG datasets. (a, c): Some geodesics obtained on UCG (a) and WCG (c), for $6$ different Riemannian metrics. The contour plots represent the energy landscape given by $-\log p_{\mathcal{M}}$. (b, d) Quantitative evaluation of geodesics on UCG (b) and WCG (d). We report (i) the accumulated probability along the geodesic (the higher the better) and ii) RMSE between each geodesic and its corresponding baseline (the lower the better). See Supp. D.3 for the $2$-$\sigma$ error.
@@
To test how different metrics behave when the density varies along the data manifold, we switch from the uniformly populated UCG semicircle to the Weighted Circular Gaussian (WCG), whose samples cluster near the arc’s centre. As shown in Fig. 3, log-based metrics (${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$, $\boldsymbol{G}_{E_{\mathcal{M}}}$) accentuate the manifold curvature more than $1/p$-based ones (${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$, ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$, ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$, $\boldsymbol{G}_{1/p_{\mathcal{M}}}$), producing larger steps in high-density regions. This is because $-\log p$ diverges as $p\to 0$, amplifying distortions and speed variations.
@@ 4.3 Rotated Characters
We use an image dataset of seven rotated, non-symmetric characters in two variants: Uniform Rotated Characters (URC), with evenly distributed angles, and Biased Rotated Characters (BRC), concentrated near $0^{\circ}$. In this subsection, all computations are done in the $64$-dimensional latent space of a regularized autoencoder trained with a triplet loss, ensuring that small angular differences yield short latent distances. This setup lets us test Riemannian metrics in high-dimensional space by treating the smooth rotation that maps one character pose to another as the ground-truth geodesic. Separate EBMs and interpolant networks are trained for each dataset variant. Full experimental details (datasets, architectures, and hyperparameters) are provided in Supp. E.
Metric $\mathcal{D}$-RMSE ()($\downarrow$)( ↓ ) $\gamma^{\star}$-RMSE ()($\downarrow$)( ↓ )
linear interp. 2.96 3.52
[…]
@@
[…]
@@
Figure 4: Geodesics on the URC dataset. (a) Geodesics computed with different Riemannian metrics, projected into pixel space for visualization. ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$ and ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ often deviate from the intended path, sometimes drifting toward other characters (e.g., the letter F). (b) Quantitative evaluation using two metrics: (i) $\mathcal{D}$-RMSE, which measures proximity to the dataset manifold, and (ii) $\gamma$-RMSE, which measures the deviation from an ideal smooth rotation. See Supp. E.6 for the $2$-$\sigma$ error.
In Fig. 4a, we visualize geodesics computed on the URC dataset, projected back into pixel space (see Supp. E.5 for additional results on both URC and BRC). EBM-based metrics (${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$ and ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$) yield smooth rotations that preserve character identity, while ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$ and especially ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ often deviate from the intended trajectory. To illustrate these failures, Fig. 1 shows all geodesics projected into PCA
@@
space for a case involving the letter F. While ${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$ and ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$ remain on the manifold of rotated F instances, linear interpolation cuts through low-density regions, and ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$ and ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ drift toward other character classes. To quantify this, we use two metrics: $\mathcal{D}$-RMSE, which measures the average distance from each geodesic point to its nearest neighbor in the dataset—lower values indicate better adherence to the data manifold; and $\gamma$-RMSE, which evaluates how closely the geodesic follows an ideal smooth rotation between endpoints. All results are averaged over $1,000$ geodesics with random endpoint orientations. As shown in Fig. 4b, EBM-based metrics consistently outperform others; ${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$ performs reasonably well, while ${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$ shows large deviations on both metrics. Overall, these results suggest that EBM-based metrics scale more effectively to high-dimensional data than alternative approaches.

### v3

@@ 1 Introduction
In this article, we answer affirmatively by proposing to derive Riemannian metrics from pretrained Energy-Based Models (EBMs) [16, 19, 20]. EBMs are a flexible class of generative models that define an energy function $E_{\theta}$, parameterized by a neural network, assigning low energy to likely data points (i.e., $p_{\theta}(\mathbf{x})\propto\exp(-E_{\theta}(\mathbf{x}))$).We show that the energy landscape of an EBM encodes a rich geometric structure and can be leveraged to derive effective Riemannian metrics. Specifically, we introduce two novel conformal Riemannian metrics—metrics that scale the identity by a positive scalar function: $\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} proportional to the energy itself, and $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}, proportional to the inverse unnormalized density. We evaluate both against established alternatives ($\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$} [13] and $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} [11]) across datasets of increasing complexity—from toy distributions with known geodesics (see § 4.2), to rotated character images where the manifold structure is partially known (see § 4.3), and finally to high-dimensional natural images where no ground truth geometry is available (see § 4.4). Throughout this work, we adopt the common choice of equipping the data space with a density-based Riemannian metric, thereby defining the geometry of the manifold in terms of data concentration. We show that EBM-based metrics yield geodesics that (i) remain closer to the data manifold and (ii) better reflect its intrinsic curvature (see Fig. 1 for a visualization of the geodesics).
@@ 3.1 A primer on Riemannian geometry
$L(\boldsymbol{\gamma})=\int_{0}^{1}\sqrt{\langle\dot{\boldsymbol{\gamma}}(t),\dot{\boldsymbol{\gamma}}(t)\rangle_{\boldsymbol{\gamma}(t)}}dt.$ (1)
In Eq. 1, $\dot{\boldsymbol{\gamma}}(t)$ denotes the velocity vector of the curve $\boldsymbol{\gamma}(t)$, which lies in the tangent space at that point (i.e., $\dot{\boldsymbol{\gamma}}(t)\in\mathcal{T}_{{\boldsymbol{\gamma}}(t)}^{\mathcal{M}}$). The minimizer of Eq. 1 is called a geodesic; it represents the (locally) shortest path between $\mathbf{x_{0}}$ and $\mathbf{x_{1}}$. In this work, we minimize the kinetic energy functional instead of the length (see Eq. 2). Although both functionals yield the same geodesics up to a parametrization, minimizing the kinetic energy functional results in a constant Riemannian speed parametrization [^1] With length fixed, the strictly convex energy $E=\tfrac{1}{2}\int_{0}^{1}v(t)^{2}dt$ attains its minimum—by Jensen’s inequality—only when the speed $v(t)$ is constant.. This property simplifies optimization and improves numerical stability [9, 13].
$\boldsymbol{\gamma}^{\star}(t)=\operatorname*{arg\,min}_{\boldsymbol{\gamma}}\mathcal{E}[\boldsymbol{\gamma}]\text{\;\;s.t.\;\;}\mathcal{E}[\boldsymbol{\gamma}]=\frac{1}{2}\int_{0}^{1}\langle\dot{\boldsymbol{\gamma}}(t),\dot{\boldsymbol{\gamma}}(t)\rangle_{\boldsymbol{\gamma}(t)}\,dt.$ (2)
@@
This parameterization satisfies the boundary conditions ($\mathbf{x}_{0,\eta}{=}\mathbf{x}_{0}$, $\mathbf{x}_{1,\eta}{=}\mathbf{x}_{1}$). In Eq. 3, $\boldsymbol{\varphi}_{\eta}$ serves as a nonlinear correction to the linear path, allowing the learned path to bend toward the data manifold. We train a single interpolant network $\boldsymbol{\varphi}_{\eta}$ over batches of random endpoint pairs so it can approximate geodesics between arbitrary points (see Algo. 1). Intuitively, our geodesic interpolant begins with a straight line between the endpoints and uses a neural network to compute a smooth curvature relative to this baseline—bending the path toward regions of higher data density, much like pulling a string taut over a curved surface that reflects the geometry of the data. Unlike Kapusniak et al. [24], who use full autodifferentiation to compute $\dot{\mathbf{x}}_{t,\eta}$, we opt for finite difference instead. We found this approach more stable and accurate when using a fine-time discretization.
@@ 3.2 Energy-Based Models
Let $p_{\mathcal{M}}$ be the true data distribution supported on the manifold $\mathcal{M}$, such that $\int_{\mathbf{x}\in\mathcal{M}}p_{\mathcal{M}}(\mathbf{x})d\mathbf{x}=1$. In practice, we do not have access to $p_{\mathcal{M}}$ directly, but only to a finite set of samples $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$drawn from it. These samples define the empirical distribution $p_{\mathcal{D}}$, which we use to train our models.
@@
$p_{\theta}(\mathbf{x})=\frac{\exp\big(-E_{\theta}(\mathbf{x})\big)}{Z(\theta)}\,\,\,\,\textnormal{where}\,\,\,Z(\theta)=\int\exp\big(-E_{\theta}(\mathbf{x})\big)d\mathbf{x}.$ (4)
@@
EBM can be hard to train in high-dimensional pixel space, especially because of the sampling procedure [53, 54, 55]. For complex tasks, we follow standard practice and operate in the latent space of a pretrained VAE [56], where all baselines are evaluated for fairness. To improve the EBM training training stability, we regularize the contrastive divergence loss with a denoising term, which preserves the global structure of the energy landscape while enhancing convergence—a technique we find both effective and broadly applicable.
@@ 3.3 EBM-derived Riemannian Metrics
Here, we describe the EBM-derived metrics $\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$}, $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}. For details on the baseline Riemannian metrics $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$$\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$}, see § 4.1. To ensure a fair comparison —and following standard practice in the field [42, 57, 58]— all metrics are cast using a shared parametric form:
@@
where $\mathbf{h}(\mathbf{x})$ is a metric-specific, positive-definite function (either scalar, diagonal, or matrix), and $\alpha$, $\beta$ are calibration constants. These constants are chosen so that the metric scale to $\mathbf{I}$ on the data manifold and to $10^{3}\cdot\mathbf{I}$ in low-density regions [^2] Note that this multiplicative factor amounts to a change of unit, to ensure reasonable scaling of the lengths, but the induced geodesics are only determined by the ratio $\alpha/\beta$.. This allows fair comparison across metric choices without introducing significant sensitivity to hyperparameter tuning. Further details about the metric calibration procedure are provided in Supp. C.1. Importantly, all EBM-derived metrics are conformal, they take the form $\lambda(\mathbf{x})\mathbf{I}$, where $\lambda$ is a scalar function. In other words, they scale the identity matrix uniformly in all directions, resulting in isotropic metrics:
@@
    Intuitively, high-energy (low-density) regions receive a larger metric, penalizing movement away from the data. Note that $E_{\theta}$ is an affine rescaling of the negative log-likelihood $\log p_{\mathcal{D}}$.
@@ 4.1 Baseline Riemannian Metrics
[…]
@@
The scaling constants ($\alpha$, $\beta$) are introduced to ensure consistent dynamic range across metrics and have minimal impact on convergence or geodesic quality; the number of discretization steps ($T=100$) is chosen as a trade-off between efficiency and accuracy, consistent with prior work. We evaluate $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}, $\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$}, $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$}, and $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} on three datasets of increasing complexity. Circular Mixture of Gaussians offers full control and ground-truth geodesics. The rotated characters dataset is higher-dimensional but still allows quantitative evaluation. Animal Faces is made of higher-dimensional images but with no ground truth. This progression tests metric performance as data complexity grows. The code to reproduce all our experiments is available at https://github.com/VictorBoutin/RiemannEBM.
@@ 4.2 Circular Mixture of Gaussians
Metric
$p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$
$(\uparrow)$
RMSE
$(\downarrow)$
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.79$ -
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.77$ -
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} $\mathbf{0.78}$ $0.12$
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} $0.60$ $0.38$
Metric
$p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$
$(\uparrow)$
RMSE
$(\downarrow)$
$\boldsymbol{G}_{E_{\mathcal{M}}}$ 0.67 -
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ 0.73 -
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} 0.67 0.18
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} 0.65 0.34
Figure 2: Geodesics on UCG and WCG datasets. (a, c): Some geodesics obtained on UCG (a) and WCG (c), for 66 different Riemannian metrics. The contour plots represent the energy landscape given by $\log p_{\mathcal{M}}$. (b, d) Quantitative evaluation of geodesics on UCG (b) and WCG (d). We report (i) the accumulated probability along the geodesic (the higher the better) and ii) RMSE between each geodesic and its corresponding baseline (i.e., $\boldsymbol{G}_{E_{\mathcal{M}}}$ for $\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$}, and $\boldsymbol{G}_{1/p_{\mathcal{M}}}$ for $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}, $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} and $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$}). See Supp. D.3 for the 22-$\sigma$ error.
@@
To test how different metrics behave when the density varies along the data manifold, we switch
@@
from the uniformly populated UCG semicircle to the Weighted Circular Gaussian (WCG), whose samples cluster near the arc’s centre. As shown in Fig. 3, log-based metrics ($\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$}, $\boldsymbol{G}_{E_{\mathcal{M}}}$) accentuate the manifold curvature more than $1/p$-based ones ($\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}, $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$}, $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$}, $\boldsymbol{G}_{1/p_{\mathcal{M}}}$), producing larger steps in high-density regions. This is because $\log p$ diverges as $p\to 0$, amplifying distortions and speed variations.
@@ 4.3 Rotated Characters
Metric
$\mathcal{D}$-RMSE
$(\downarrow)$
$\gamma^{\star}$-RMSE
$(\downarrow)$
linear
interp.
2.96 3.52
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} 0.11 0.40
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} 0.66 2.39
@@
Figure 4: Geodesics on the URC dataset. (a) Geodesics computed with different Riemannian metrics, projected into pixel space for visualization. $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$} and $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} often deviate from the intended path, sometimes drifting toward other characters (e.g., the letter F). (b) Quantitative evaluation using two metrics: (i) $\mathcal{D}$-RMSE, which measures proximity to the dataset manifold, and (ii) $\gamma$-RMSE, which measures the deviation from an ideal smooth rotation. See Supp. E.6 for the 22-$\sigma$ error.
We use an image dataset of seven rotated, non-symmetric characters in two variants: Uniform Rotated Characters (URC), with evenly distributed angles, and Biased Rotated Characters (BRC), concentrated near $0^{\circ}$. In this subsection, all computations are done in the $64$-dimensional latent space of a regularized autoencoder trained with a triplet loss, ensuring that small angular differences yield short latent distances. This setup provides a unique middle ground: although the underlying Riemannian metric is unknown, we can treat the smooth in-plane rotation between two instances of the same character as a proxy for the ground-truth geodesic. Thanks to the triplet loss, the latent space is structured so that nearby points correspond to slight rotations of the same character, making the shortest path between two orientations a meaningful approximation of the true geodesic in the task-relevant transformation space. Separate EBMs and interpolant networks are trained for each dataset variant. Full experimental details (datasets, architectures, and hyperparameters) are provided in Supp. E.
In Fig. 4a, we visualize geodesics computed on the URC dataset, projected back into
@@
pixel space (see Supp. E.5 for additional results on both URC and BRC). EBM-based metrics ($\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} and $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$}) yield smooth rotations that preserve character identity, while $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$} and especially $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} often deviate from the intended trajectory. To illustrate these failures, Fig. 1 shows all geodesics projected into PCA space for a case involving the letter F. While $\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} and $\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}$} remain on the manifold of rotated F instances, linear interpolation cuts through low-density regions, and $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$} and $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} drift toward other character classes. To quantify this, we use two metrics: $\mathcal{D}$-RMSE, which measures the average distance from each geodesic point to its nearest neighbor in the dataset—lower values indicate better adherence to the data manifold; and $\gamma$-RMSE, which evaluates how closely the geodesic follows an ideal smooth rotation between endpoints. All results are averaged over $1,000$ geodesics with random endpoint orientations. As shown in Fig. 4b, EBM-based metrics consistently outperform others; $\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}$} performs reasonably well, while $\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} shows large deviations on both metrics. Overall, these results suggest that EBM-based metrics scale more effectively to high-dimensional data than alternative approaches.